OpenStudy (nataliejlr):

Evaluate and simplify the expression when c=1 and d=3. 3( 2c+d ) - 4( c-d ) + d^2 = ?

OpenStudy (vitalina):

plug in all the variables: 3(2*1+3) - 4(1-3) + 3^2 now solve (following the order of operations PEMDAS): = 3(2+3) - 4(-2) + 9 = 3(5) - (-8) + 9 = 15 + 8 + 9 = 32